# Building Symbiotic AI: Reviewing the AI Act for a Human-Centred, Principle-Based Framework

### v1

@@ Abstract
Artificial Intelligence (AI) spreads quickly as new technologies and services take over modern society. The need to regulate AI design, development, and use is strictly necessary to avoid unethical and potentially dangerous consequences to humans. The European Union (EU) has released a new legal framework, the AI Act, to regulate AI by undertaking a risk-based approach to safeguard humans during interaction. At the same time, researchers offer a new perspective on AI systems, commonly known as Human-Centred AI (HCAI), highlighting the need for a human-centred approach to their design. In this context, Symbiotic AI (a subtype of HCAI) promises to enhance human capabilities through a deeper and continuous collaboration between human intelligence and AI. This article presents the results of a Systematic Literature Review (SLR) that aims to identify principles that characterise the design and development of Symbiotic AI systems while considering humans as the core of the process. Through content analysis, four principles emerged from the review that must be applied to create Human-Centred AI systems that can establish a symbiotic relationship with humans. In addition, current trends and challenges were defined to indicate open questions that may guide future research for the development of SAI systems that comply with the AI Act.
@@ Introduction
These concerns have prompted academics and global legislative bodies to take action to ensure the responsible development of AI. In the academic sphere, a new perspective known as Human-Centred Artificial Intelligence (HCAI) [5, 6] is reshaping the research landscape. Located at the intersection between Human-Computer Interaction (HCI) and AI, HCAI promises to provide a direction to design AI systems that are safe, reliable, and trustworthy [7]. More precisely, HCAI aims to design, develop, and evaluate AI systems involving end-users in the process to increase their performances and satisfaction in performing specified tasks [8]. Therefore, HCAI systems aim to be usable and useful for specified users to reach their specified goals in their context of use, while being reliable, safe to use, and trustworthy [8]. By adopting a human-centred design approach, HCAI may foster a symbiotic relationship between humans and AI. Therefore, Symbiotic Artificial Intelligence (SAI) systems are a specific type of HCAI systems that allow for continuing and deeper collaborations between human intelligence and AI, aiming at their mutual augmentations without hampering human autonomy or AI’s performance [9, 8]. However, this may not suffice to ease ethical concerns.
@@
Simultaneously, governments are beginning to draft and implement new legislative measures. A notable example is the European Union (EU)’s AI Act, which represents a pioneering effort in regulating AI [10]. The AI Act is a legal framework concerning AI systems’ design, development, deployment, and use, aiming at ensuring their proper employment while minimising risks to human well-being and society. It leverages a human-centric and risk-based approach, considering humans in all their dimensions and not as mere users, promoting fundamental rights and categorising AI-based systems into four distinct risk levels: Low-Minimal, Limited, High, and Unacceptable. This classification is based on the AI system’s purpose and its potential impact on human well-being, with stricter rules applied as the level of risk increases [10].
@@
This work aims to identify the principles that can guide the design of SAI systems to comply with the current legal landscape considering humans as the core of the development process [11]. With this goal in mind, we performed a Systematic Literature Review (SLR), employing content analysis to understand what researchers suggest about the relationship between SAI (and generally HCAI) and the AI Act. The main contribution of this study is the proposal of principles to guide stakeholders in the design of SAI systems, i.e., of HCAI systems that favour a symbiotic relationship with humans. In addition, three properties strictly connected to the principles emerged that are represented in a value-based framework shown in Figure 2. It is finally put in the context of the relevant literature, comparing it with other principled-based approaches to SAI and HCAI. The goal of this study is thus better formulated in the following research question:
@@
What principles can guide the design of SAI systems, while promoting a symbiotic human-AI relationship, and adhering to the European AI Act?
@@
The article is structured as follows: Section ‘Results: A Principled Human-Centred AI Framework’ presents the results obtained by the SLR (i.e., principles and properties that characterize the human-AI symbiosis); Section ‘Discussion’ discusses the analysis of the literature presenting the challenges, trends, and limitations faced while conducting the SLR; Section ‘Methods’ describes the method adopted to conduct the SLR.
Results: A Principled Human-Centred AI Framework
@@
Principles for Human-AI Symbiosis
@@
Explainability aims to provide explanations concerning AI systems’ operations, ensuring that even when humans cannot understand how an AI system provides an output, they can at least receive information about why it was produced [18]. In other words, this aspect contributes to the degree to which an AI system is open and observable to humans [14, 16]. When interacting with AI systems, humans should have the right to explanations to make informed decisions, supported by AI and not replaced by it, [19, 20], while being able to oversee its processes [21].
@@
Properties of Human-Centred AI
Three properties emerged from the SLR, which underlie the principles that were identified and that must characterise AI-based systems that are designed and developed compliantly with the AI Act. The analysis was performed using hierarchical clustering, and after the initial step of identifying the first set of clusters, a hierarchical grouping was brought to the identification of the principles described in the previous section. A further grouping step allowed us to identify the overall properties of a truly human-AI symbiotic relationship: Trustworthiness, Robustness, and Sustainability. This section details these properties, highlighting their relationship with the four principles described in Section ‘Results: A Principled Human-Centred AI Framework’. It is highlighted that these properties are not a sufficient condition for AI systems to be symbiotic, but they emerged as crucial components of the design phase from this SLR, influenced by its goals. Figure 2 shows the connection among the identified properties and principles.
@@ Discussion
Grounding the Framework in the Literature
To better ground the framework introduced in Section ‘Properties of Human-Centred AI’, this section compares it with existing similar frameworks available in the literature. More precisely, three sets of principles, are taken as reference for the comparison [33, 42, 6]. The comparison among them and the proposed human-centered SAI principles is shown in Figure 3.
@@
Uncovering Current Trends and Challenges
@@
Methods
  1. (Q1)
    european AND (“artificial intelligence” OR ai) AND act
  2. (Q2)
  3. (Q3)
  4. (Q4)
  5. (Q5)
Inclusion and Exclusion Criteria
  • Language: each paper not written in English was excluded.
Coding and Classification
Conducting the SLR
Figure 5: PRISMA diagram of the literature review
The identified 58 publications are presented in LABEL:tab:map-articles-principles, showing the mapping with the principles and their dimensions. Table 2 provides an overview of which article discusses a property in the context of each principle. The framework represented in Figure 2 is elicited through a qualitative analysis of the data represented in LABEL:tab:map-articles-principles and 2.
Appendix A Classification of the Selected Manuscripts
@@
LABEL:tab:map-articles-principles and 2 present the classification of the selected articles, providing two different views. LABEL:tab:map-articles-principles presents, for each selected article, the various dimensions that are associated with all the principles. Table 2 provides an overview of which article discusses a property in the context of each principle.
Table 1: Mapping between the articles and the principles
Ref. Protection (P1) Transparency (P2) Fairness (P3) Automation Level (P4)
[12]
Safety
[47]
Interpretability
Human-In-The-Loop;
Human-On-The-Loop
[19]
Explainability
[48]
Explainability
Human-On-The-Loop
[16]
Explainability;
Interpretability
Human-On-The-Loop
[31]
Security
Explainability
Human-In-The-Loop
[21]
Safety
Human-On-The-Loop
[27]
Human-On-The-Loop
[34]
Safety
[15]
Explainability
[49]
Privacy
Non-Discrimination
[50]
Privacy;
Security
[41]
Safety
Human-On-The-Loop
[51]
Rightful Information
[52]
Human-In-The-Loop
[30]
Privacy;
Safety;
Security
Non-Discrimination
Human-In-The-Loop;
Human-On-The-Loop
[53]
Safety
Human-In-The-Loop
[22]
Privacy
[54]
Privacy;
Security
[39]
Safety
[55]
Privacy
[56]
Explainability
Human-On-The-Loop
[57]
Human-In-The-Loop
[23]
Rightful Information
Human-On-The-Loop
[58]
Privacy;
Safety
[24]
Rightful Information
[59]
Privacy
[25]
Human-On-The-Loop
[29]
Human-In-The-Loop;
Human-On-The-Loop
[35]
Explainability;
Interpretability
Non-Discrimination
Human-In-The-Loop;
Human-On-The-Loop
[60]
Privacy;
Safety;
Security
[61]
Privacy;
Safety
[62]
Safety;
Security
Human-On-The-Loop
[63]
Safety
[64]
Safety
[36]
Explainability
Non-Discrimination
Human-In-The-Loop
[65]
Privacy
Rightful Information
[26]
Privacy
[66]
Explainability;
Interpretability
[14]
Privacy;
Security
Explainability;
Interpretability
Human-On-The-Loop
[18]
Security
Explainability
Rightful Information
Human-In-The-Loop;
Human-On-The-Loop
[38]
Safety;
Security
Explainability
Human-On-The-Loop
[20]
Safety;
Security
Explainability
[13]
Privacy;
Safety
Non-Discrimination;
Rightful Information
Human-On-The-Loop
[67]
Privacy
Explainability
Non-Discrimination
Human-In-The-Loop
[68]
Privacy;
Safety
Human-On-The-Loop
[69]
Privacy;
Safety
Non-Discrimination;
Rightful Information
[70]
Safety
Non-Discrimination;
Rightful Information
Human-In-The-Loop;
Human-On-The-Loop
[71]
Privacy;
Security
Explainability;
Interpretability
Human-In-The-Loop
[72]
Safety
Human-On-The-Loop
[73]
Safety
Human-On-The-Loop
[74]
Explainability;
Interpretability
Human-On-The-Loop
[75]
Explainability
Human-In-The-Loop;
Human-On-The-Loop
[76]
Privacy
Explainability
[77]
Privacy;
Safety
[78]
Privacy;
Safety
Non-Discrimination;
Rightful Information
[79]
Explainability;
Interpretability
Non-Discrimination
Human-In-The-Loop
[80]
Safety
Non-Discrimination
Human-On-The-Loop
Table 1: (continued)
Table 2: Mapping the articles between principles and properties
Trustworthiness Robustness Sustainability
Protection [31, 21, 34, 49, 41, 30, 53, 22, 54, 39, 60, 62, 64, 26, 18, 20, 13, 67, 68, 69, 70, 71, 73, 74, 75, 76, 78, 79, 80] [31, 30, 54, 26, 14, 20, 71, 73, 38, 77]
Transparency [16, 14, 18, 20, 67, 71, 74, 75, 76, 79, 47, 48, 15, 56, 35, 69, 51, 23, 24] [31, 18, 35, 39]
Fairness [49, 18, 13, 67, 69, 70, 78, 79, 80, 35, 36] [18, 13, 69]
Automation Level [16, 48, 31, 21, 41, 30, 53, 62, 18, 13, 67, 68, 70, 71, 73, 74, 75, 80, 56, 35, 69, 23, 27, 25, 29, 12]

### v2

# Building Symbiotic Artificial Intelligence: Reviewing the AI Act for a Human-Centred, Principle-Based Framework

@@ Abstract
Artificial Intelligence (AI) spreads quickly as new technologies and services take over modern society. The need to regulate AI design, development, and use is strictly necessary to avoid unethical and potentially dangerous consequences to humans. The European Union (EU) has released a new legal framework, the AI Act, to regulate AI by undertaking a risk-based approach to safeguard humans during interaction. At the same time, researchers offer a new perspective on AI systems, commonly known as Human-Centred AI (HCAI), highlighting the need for a human-centred approach to their design. In this context, Symbiotic AI (a subtype of HCAI) promises to enhance human capabilities through a deeper and continuous collaboration between human intelligence and AI. This article presents the results of a Systematic Literature Review (SLR) that aims to identify principles that characterise the design and development of Symbiotic AI systems while considering humans as the core of the process. Through content analysis, we elicit four principles that must be applied to create Human-Centred AI systems that can establish a symbiotic relationship with humans. In addition, current trends and challenges are presented to indicate open questions that may guide future research for the development of SAI systems that comply with the AI Act.
keywords:
Systematic Literature Review, Artificial Intelligence, Human-Centred, AI Act
@@ 1 Introduction
These concerns have prompted academics and global legislative bodies to take action to ensure the responsible development of AI. In the academic sphere, a new perspective known as Human-Centred Artificial Intelligence (HCAI) [80, 66] is reshaping the research landscape. Located at the intersection between Human-Computer Interaction (HCI) and AI, HCAI promises to provide a direction to design AI systems that are safe, reliable, and trustworthy [65]. More precisely, HCAI aims to design, develop, and evaluate AI systems involving end-users in the process to increase their performances and satisfaction in performing specified tasks [11]. Therefore, HCAI systems aim to be usable and useful for specified users to reach their specified goals in their context of use, while being reliable, safe to use, and trustworthy [11]. By adopting a human-centred design approach, HCAI may foster a symbiotic relationship between humans and AI. Therefore, Symbiotic Artificial Intelligence (SAI) systems are a specific type of HCAI systems that allow for continuing and deeper collaborations between human intelligence and AI, aiming at their mutual augmentations without hampering human autonomy or AI’s performance [17, 11, 72]. However, this may not suffice to ease ethical concerns.
@@
Simultaneously, governments are beginning to draft and implement new legislative measures. A notable example is the European Union (EU)’s AI Act, which represents a pioneering effort in regulating AI [14]. The AI Act is a legal framework concerning AI systems’ design, development, deployment, and use, aiming at ensuring their proper employment while minimising risks to human well-being and society. It leverages a human-centric and risk-based approach [44], considering humans in all their dimensions and not as mere users, promoting fundamental rights and categorising AI-based systems into four distinct risk levels: Low-Minimal, Limited, High, and Unacceptable. This classification is based on the AI system’s purpose and its potential impact on human well-being, with stricter rules applied as the level of risk increases [14].
@@
This work aims to identify the principles that can guide the design of SAI systems to comply with the current legal landscape considering humans as the core of the development process [29]. With this goal in mind, we performed a Systematic Literature Review (SLR), employing content analysis to understand what researchers suggest about the relationship between SAI and the AI Act. The main contribution of this study is the proposal of principles to guide stakeholders in the design of SAI systems, i.e., of HCAI systems that favour a symbiotic relationship with humans. In addition, three properties strictly connected to the principles emerged that are represented in a value-based framework shown in Figure 3. It is finally put in the context of the relevant literature, comparing it with other principled-based approaches to SAI and HCAI. The goal of this study is thus better formulated in the following research question:
@@
RQ: What principles can guide the design of SAI systems, while promoting a symbiotic human-AI relationship, and adhering to the European AI Act?
@@
The article is structured as follows: Section 2 describes the method adopted to conduct the SLR; Section 3 presents the results obtained by the SLR (i.e., principles and properties that characterize the human-AI symbiosis); Section 4 discusses the analysis of the literature presenting the challenges, trends, and limitations faced while conducting the SLR.
2 Methods
  1. (Q1)
    european AND (“artificial intelligence” OR ai) AND act
  2. (Q2)
  3. (Q3)
  4. (Q4)
  5. (Q5)
2.1 Inclusion and Exclusion Criteria
  • Language: each paper not written in English was excluded.
2.2 Coding and Classification
2.3 Conducting the SLR
Figure 1: PRISMA diagram of the literature review
The identified 58 publications are presented in LABEL:tab:map-articles-principles, showing the mapping with the principles and their dimensions. LABEL:tab:map-principles-properties provides an overview of which article discusses a property in the context of each principle. The framework represented in Figure 3 is elicited through a qualitative analysis of the data represented in LABEL:tab:map-articles-principles and LABEL:tab:map-principles-properties.
3 Results: A Principled Human-Centred AI Framework
@@
3.1 Principles for Human-AI Symbiosis
@@ 3.1.1 Transparency
Explainability aims to provide explanations concerning AI systems’ operations, ensuring that even when humans cannot understand how an AI system provides an output, they can at least receive information about why it was produced [67]. In other words, this aspect contributes to the degree to which an AI system is open and observable to humans [58, 75]. When interacting with AI systems, humans should have the right to explanations to make informed decisions, supported by AI and not replaced by it [39, 78], while being able to oversee its processes [74].
@@ 3.1.4 Protection
3.2 Properties of Human-Centred AI
Three properties emerged from the SLR, which underlie the principles that were identified and that must characterise AI-based systems that are designed and developed compliantly with the AI Act. The analysis was performed using hierarchical clustering, and after the initial step of identifying the first set of clusters, a hierarchical grouping was brought to the identification of the principles described in the previous section. A further grouping step allowed us to identify the overall properties of a truly human-AI symbiotic relationship: Trustworthiness, Robustness, and Sustainability. This section details these properties, highlighting their relationship with the four principles described in Section 3. It is highlighted that these properties are not a sufficient condition for AI systems to be symbiotic, but they emerged as crucial components of the design phase from this SLR, influenced by its goals. Figure 3 shows the connection among the identified properties and principles.
@@ 4 Discussion
4.1 Grounding the Framework in the Literature
To better ground the framework introduced in Section 3.2, this section compares it with existing similar frameworks available in the literature. More precisely, three sets of principles, are taken as reference for the comparison [15, 41, 66]. The comparison among them and the proposed human-centered SAI principles is shown in Figure 4.
@@ 4.1 Grounding the Framework in the Literature
4.2 Uncovering Current Trends and Challenges
@@ 4.3 Limitations
Appendix A Appendix
A.1 Classification of the Selected Manuscripts
@@
LABEL:tab:map-articles-principles and LABEL:tab:map-principles-properties present the classification of the selected articles, providing two different views. LABEL:tab:map-articles-principles presents, for each selected article, the various dimensions that are associated with all the principles. LABEL:tab:map-principles-properties provides an overview of which article discusses a property in the context of each principle.
{longtblr}
[ caption = Mapping between the articles and the principles, label = tab:map-articles-principles ] colspec = m3cmccccc, rowhead = 1, rows = font= Ref. Protection (P1) Transparency (P2) Fairness (P3) Automation Level (P4)
[Helberger2023ChatGPT] Safety
[Lazcoz2023Humans] Interpretability Human-In-The-Loop; Human-On-The-Loop
[Malgieri2024Licensing] Explainability
[Varosanec2022Path] Explainability Human-On-The-Loop
[Urquhart2022Legal] Explainability; Interpretability Human-On-The-Loop
[Kieseberg2023Controllable] Security Explainability Human-In-The-Loop
[Tartaro2023European] Safety Human-On-The-Loop
[Enqvist2023Human] Human-On-The-Loop
[Neuwirth2023Prohibited] Safety
[Gyevnar2023Bridging] Explainability
[Stuurman2022Regulating] Privacy Non-Discrimination
[Mueck2023Upcoming] Privacy; Security
[Hacker2023European] Safety Human-On-The-Loop
[Porlezza2023Promoting] Rightful Information
[Chamberlain2023RiskBased] Human-In-The-Loop
[Hupont2022Landscape] Privacy; Safety; Security Non-Discrimination Human-In-The-Loop; Human-On-The-Loop
[Mylly2023Transparent] Safety Human-In-The-Loop
[Anamaria2023Artificial] Privacy
[Hupont2023Documenting] Privacy; Security
[Pagallo2022Environmental] Safety
[Morozovaite2023Hypernudging] Privacy
[Laux2023Institutionalised] Explainability Human-On-The-Loop
[Mazur2024Embedding] Human-In-The-Loop
[Herbosch2024Fraud] Rightful Information Human-On-The-Loop
[RomeroMoreno2024Generative] Privacy; Safety
[Sposini2024Neuromarketing] Rightful Information
[Yordanova2024Regulating] Privacy
[Biewer2024Software] Human-On-The-Loop
[CoveloDeAbreu2024Artificial] Human-In-The-Loop; Human-On-The-Loop
[Stettinger2024Trustworthiness] Explainability; Interpretability Non-Discrimination Human-In-The-Loop; Human-On-The-Loop
[Almada2024Brussels] Privacy; Safety; Security
[Aseeva2023Liable] Privacy; Safety
[Helberger2024FutureNewsCorp] Safety; Security Human-On-The-Loop
[Hirvonen2023Just] Safety
[Hupont2024Use] Safety
[Laux2024Trustworthy] Explainability Non-Discrimination Human-In-The-Loop
[Migliorini2024More] Privacy Rightful Information
[Nikiforov2024Groups] Privacy
[Olsen2024Right] Explainability; Interpretability
[Pavlidis2024Unlocking] Privacy; Security Explainability; Interpretability Human-On-The-Loop
[Siegel2024Media] Security Explainability Rightful Information Human-In-The-Loop; Human-On-The-Loop
[Tallberg2024AI] Safety; Security Explainability Human-On-The-Loop
[Wagner2024Navigating] Safety; Security Explainability
[Worsdorfer2023Mitigating] Privacy; Safety Non-Discrimination; Rightful Information Human-On-The-Loop
[Kattnig2024Assessing] Privacy Explainability Non-Discrimination Human-In-The-Loop
[Montagnani2024EU] Privacy; Safety Human-On-The-Loop
[Laux2024Three] Privacy; Safety Non-Discrimination; Rightful Information
[BoteroArcila2024AI] Safety Non-Discrimination; Rightful Information Human-In-The-Loop; Human-On-The-Loop
[Baumgartner2024AIDriven] Privacy; Security Explainability; Interpretability Human-In-The-Loop
[VanKolfschooten2024EU] Safety Human-On-The-Loop
[Lee2024Transformative] Safety Human-On-The-Loop
[CanteroGamito2024Artificial] Explainability; Interpretability Human-On-The-Loop
[Peeters2024Editorial] Explainability Human-In-The-Loop; Human-On-The-Loop
[Nisevic2024Explainable] Privacy Explainability
[Novelli2024Robust] Privacy; Safety
[Cupac2024Regulate] Privacy; Safety Non-Discrimination; Rightful Information
[Papadakis2024Explainable] Explainability; Interpretability Non-Discrimination Human-In-The-Loop
[Gornet2024European] Safety Non-Discrimination Human-On-The-Loop
{longtblr}
[ caption = Mapping the articles between principles and properties, label = tab:map-principles-properties ] colspec = m2cmXXX, rowhead = 1, rows = font= & Trustworthiness Robustness Sustainability
Protection [Kieseberg2023Controllable, Tartaro2023European, Neuwirth2023Prohibited, Stuurman2022Regulating, Hacker2023European, Hupont2022Landscape, Mylly2023Transparent, Anamaria2023Artificial, Hupont2023Documenting, Pagallo2022Environmental, Almada2024Brussels, Helberger2024FutureNewsCorp, Hupont2024Use, Nikiforov2024Groups, Siegel2024Media, Wagner2024Navigating, Worsdorfer2023Mitigating, Kattnig2024Assessing, Montagnani2024EU, Laux2024Three, BoteroArcila2024AI, Baumgartner2024AIDriven, Lee2024Transformative, CanteroGamito2024Artificial, Peeters2024Editorial, Nisevic2024Explainable, Cupac2024Regulate, Papadakis2024Explainable, Gornet2024European] [Kieseberg2023Controllable, Hupont2022Landscape, Hupont2023Documenting, Nikiforov2024Groups, Pavlidis2024Unlocking, Wagner2024Navigating, Baumgartner2024AIDriven, Lee2024Transformative, Tallberg2024AI, Novelli2024Robust]
Fairness [Stuurman2022Regulating, Siegel2024Media, Worsdorfer2023Mitigating, Kattnig2024Assessing, Laux2024Three, BoteroArcila2024AI, Cupac2024Regulate, Papadakis2024Explainable, Gornet2024European, Stettinger2024Trustworthiness, Laux2024Trustworthy] [Siegel2024Media, Worsdorfer2023Mitigating, Laux2024Three]
Automation Level [Urquhart2022Legal, Varosanec2022Path, Kieseberg2023Controllable, Tartaro2023European, Hacker2023European, Hupont2022Landscape, Mylly2023Transparent, Helberger2024FutureNewsCorp, Siegel2024Media, Worsdorfer2023Mitigating, Kattnig2024Assessing, Montagnani2024EU, BoteroArcila2024AI, Baumgartner2024AIDriven, Lee2024Transformative, CanteroGamito2024Artificial, Peeters2024Editorial, Gornet2024European, Laux2023Institutionalised, Stettinger2024Trustworthiness, Laux2024Three, Herbosch2024Fraud, Enqvist2023Human, Biewer2024Software, CoveloDeAbreu2024Artificial, Helberger2023ChatGPT]